OpenStudy (anonymous):

I need help with the chain rule and implicit differentiation. e^ycosx=1+sinxy

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

I am having a really hard time with the chain rule, period. Here's where I'm at:

OpenStudy (anonymous):

\[\frac{ d }{ dx }(e ^{y})\times \frac{ d }{ dx }(cosx)=\frac{ d }{ dx }(1+sinxy)\]

OpenStudy (zarkon):

stop there

OpenStudy (zarkon):

remember that y is a function of x....so \[e^y\cos(x)\] is the product of two functions of x...thus you need to use the product rule

OpenStudy (anonymous):

right

OpenStudy (zarkon):

the way you have it written above is not correct then

OpenStudy (anonymous):

I'm having trouble with the equation tool on OpenStudy.

OpenStudy (anonymous):

I'm going to have to fix my computer and come back. Sorry. I didn't mean to waste your time.

OpenStudy (zarkon):

\[\frac{d}{dx}[e^y\cos(x)]=\frac{d}{dx}[e^y]\cdot\cos(x)+e^y\frac{d}{dx}\cdot[\cos(x)]\]